OpenStudy (anonymous):

solve the trigonometric equation sin 2x-sin x +2cos x -1=0 in the interval [0,2TT)

OpenStudy (alexwee123):

is this the equation? \[\sin ^{2}x-sinx+2cosx-1=0?\] or \[\sin2x-sinx+2cosx-1=0?\]

OpenStudy (dumbcow):

its most likely the sin(2x) option its easier and solutions are easily in terms of pi

OpenStudy (alexwee123):

ugh i didn't do trig identities in a while so.... \[\sin2x=2sinxcosx\] so... \[2sinxcosx+2cosx=sinx+1\]\[2cosx(sinx+1)=sinx+1\]\[2cosx=1\]\[cosx=1/2\]\[\cos ^{-1}(\cos(x))=\cos ^{-1}(1/2)\]\[x=\Pi/3\] did i make a mistake @dumbcow ? o.0

OpenStudy (anonymous):

sinx+1=1? only when squared

OpenStudy (dumbcow):

hmm i didn't see that solution but i see no flaw in your work

OpenStudy (anonymous):

wait its sin^2 + cos^2 = 1

OpenStudy (alexwee123):

ya ^^

OpenStudy (anonymous):

i wish i knew the answer to this problem its gonna bug me lol

OpenStudy (alexwee123):

@dumbcow what did you get as an answer? sorry that kind of sounded rude :/

OpenStudy (anonymous):

pie/3, 5pie/3 is what i got

OpenStudy (dumbcow):

no its fine...i went back and realized i did it wrong you are correct but there are other solutions...instead of cancelling out the sinx +1 factor it out \[\rightarrow (\sin x +1)(2\cos x -1) = 0\] \[x = \pi/3 , 3\pi/2, 5\pi/3\]

OpenStudy (alexwee123):

lol ya i see now :/

OpenStudy (alexwee123):

but 3pi/2 is extraneous

OpenStudy (anonymous):

i dont think 3pie/2 is part of the answer